Mathematics and Publishing on the Web

Jon Borwein Shrum Chair

Centre for Experimental & Constructive Mathematics Simon Fraser University jborwein@cecm.sfu.ca

I shall make an informal description of some of the central issues involved in electronic/internet publication, especially of mathematics. In light of our various experiences at CECM as a web-provider, I am ready to raise many questions but to provide few answers.